OpenStudy (anonymous):

free medal if you help me 80% of a number is 96. What is 25% of that same number? A. 20 B. 30 C. 40 D. 120

OpenStudy (anonymous):

hey is it A

OpenStudy (andrewhaze):

are you looking for 25% of it

OpenStudy (andrewhaze):

@rasheem0

OpenStudy (anonymous):

yes

OpenStudy (andrewhaze):

or 80%

OpenStudy (andrewhaze):

I think A

OpenStudy (andrewhaze):

@rasheem0

OpenStudy (andrewhaze):

could I get a medal

OpenStudy (andrewhaze):

@rasheem0 ^

OpenStudy (anwaarullah):

its B,,,96/x=80/100 80x=9600 x=9600/80 x=120 25%120= 30

OpenStudy (anwaarullah):

check 96/120*100= 80%